1. RANK the following four rules from MOST combinations to LEAST combinations. Explain how you know. (Intuition is good, but math is better!)
Password must be 8 lowercase letters. No numbers.
Password must be 6 letters but they may be lowercase or uppercase. No numbers.
Password must be 4 characters - either uppercase, lowercase, or numbers.
Password must be 10 letters, but each character must alternate between an UPPERCASE letter and a number.
2. How many ways are there to rank the options?

Answer by Edwin McCravy(20055)

1. RANK the following four rules from MOST combinations to LEAST combinations. Explain how you know. (Intuition is good, but math is better!)
Password must be 8 lowercase letters. No numbers.
There are 26 lower case letters, we choose 8, so

C(26,8) = 1562274

Password must be 6 letters but they may be lowercase or uppercase. No numbers.
There are 26 lower case letters amd 26 upper case letters.  That's 52 letters.
We choss 6, so

C(52,6) = 20358520 

Password must be 4 characters - either uppercase, lowercase, or numbers.
There are 52 upper and lower case letters plus 10 digits, so that's 62
total characters.  We choose 4, so

C(62,6) = 61474519

Password must be 10 letters, but each character must alternate between an UPPERCASE letter and a number.
Examples: P6U2P5R6H7, 8E6K9K0P4D

Passwords which begin with an upper case letter:

26*10*26*10*26*10*26*10*26*10 = 1188137600000

Passwords which begin with a digit:

10*26*10*26*10*26*10*26*10*26 = 1188137600000

That's a total of 1188137600000 + 1188137600000 = 2376275200000

2. How many ways are there to rank the options?
The way the problem lists them is from LEAST to MOST,
so reverse them.

Edwin